OpenStudy (anonymous):

what are the steps when u integrate the square root of x^3dx

OpenStudy (anonymous):

square root means half...so its half times 3 ??to get x^3/2

OpenStudy (campbell_st):

rewrite it in index form \[\sqrt{x^3} = x^{3/2}\] \[\int\limits x^{3/2} dx = 2/5 x ^{5/2} + c\]

OpenStudy (anonymous):

can u explain the last step please

OpenStudy (campbell_st):

when integrating 1st step is add 1 to the power... then its 1/power as the coefficient \[\int\limits x^n dx = 1/(n+1) x^{n + 1} + c\]

OpenStudy (anonymous):

how u got 2 at the last step?

OpenStudy (anonymous):

how did u get 2x??

OpenStudy (campbell_st):

i got \[2x^{2/5}/5\]

OpenStudy (campbell_st):

the power is 3/2 then add 1... 3/2 + 1 = 5/2 the reciprocal is 2/5

OpenStudy (anonymous):

okay i get it now..thxx:)